Problem 5A-7 Cost Behavior; High-Low Method; Contribution Format Income Statement [LO5-10]

Morrisey & Brown, Ltd., of Sydney is a merchandising company that is the sole distributor of a product that is increasing in popularity among Australian consumers. The company’s income statements for the three most recent months follow:

Morrisey & Brown, Ltd.
Income Statements
For the Three Months Ended September 30
July August September
Sales in units 7,750 8,250 8,750
Sales $ 775,000 $ 825,000 $ 875,000
Cost of goods sold 465,000 495,000 525,000
Gross margin 310,000 330,000 350,000
Selling and administrative expenses:
Advertising expense 28,200 28,200 28,200
Shipping expense 89,500 94,500 99,500
Salaries and commissions 140,000 147,500 155,000
Insurance expense 7,200 7,200 7,200
Depreciation expense 21,600 21,600 21,600
Total selling and administrative expenses 286,500 299,000 311,500
Net operating income $ 23,500 $ 31,000 $ 38,500

Required:

1. By analyzing the data from the company's income statements, classify each of its expenses (including cost of goods sold) as either variable, fixed, or mixed.

2. Using the high-low method, separate each mixed expense into variable and fixed elements. Express the variable and fixed portions of each mixed expense in the form Y = a + bX.

3. Redo the company’s income statement at the 8,750-unit level of activity using the contribution format.
Answer #1
1) Expenses Classification
cost of goods sold Variable
Advertising expense fixed
shipping expense mixed
Salaries and commission mixed
insurance expense fixed
Depreciation expense fixed
2) High low method
shipping expense
units cost
high 8,750 99,500
low 7,750 89,500
difference 1,000 10,000
variable expense per unit 10000/1000 = $10
fixed expense = 99500-8750*10             = 12000
Salaries and commission
units cost
high 8,750 155,000
low 7,750 140,000
difference 1,000 15,000
Variable expense per unit = 15000/1000=$15
fixed expense = 155000-8750*15 23750
cost formula
Variable cost Fixed cost                    Formula                                                       
shipping expense 10 per unit 12,000 Y= 12,000 + 10 per unit
Salaries & commission 15 per unit 23,750 Y= 23,750 + 15 per unit
3) Sales in units 8750
Sales 875,000
Variable expense
Cost of goods sold 525,000
Shipping expense 87,500
Salaries and commission 131250
743,750
Contribution margin 131,250
Fixed expense
Advertising expense 28,200
Shipping expense 12,000
Salaries & commission 23,750
insurance expense 7,200
Depreciation expense 21,600
92,750
Net income 38,500